OpenStudy (anonymous):

What are the period and phase shift for f(x) = –4 tan(x – π)?

OpenStudy (anonymous):

The general form of the equation is f(x) = Atan(ωx - φ) comparing with f(x) = –4 tan(x – π) w=1 Then, period is equal to: T=π/w (Because the period of tang is π, no 2π) T=π/1 T=π For the phase shift: phase shift = −π (i.e. shifted π units to left)